OpenStudy (anonymous):

use the remainder theorem to find the remainder when p(x)=x^4-9x^3-5x^2-3x+4 is divided by x+3

OpenStudy (xapproachesinfinity):

just do long division

OpenStudy (xapproachesinfinity):

that is really what the reminder thm is about

OpenStudy (xapproachesinfinity):

see here

OpenStudy (xapproachesinfinity):

http://www.mathsisfun.com/algebra/polynomials-remainder-factor.html

OpenStudy (xapproachesinfinity):

you do the long division you get remainder from that and this is ur answer

OpenStudy (anonymous):

I need help

OpenStudy (xapproachesinfinity):

you don't know how to do long division

OpenStudy (anonymous):

Ive doing this question for 3 hours

OpenStudy (mathmate):

Put x=-3 and evaluate the expression, i.e evaluate p(-3). -3 comes from solving x+3=0. That's the remainder theorem. See for example: http://www.purplemath.com/modules/remaindr.htm

OpenStudy (anonymous):

this is hard

OpenStudy (butterflydreamer):

So this is your equation p(x)=x^4-9x^3-5x^2-3x+4 . Using remainder theorem, we are dividing your polynomial by (x+3) so this means since (x+3) = 0..." x = -3". SO we have to substitute x = -3 into your equation p(x)=x^4-9x^3-5x^2-3x+4 (note: see the all the x's in your equation? just replace it with -3 and then plug it into the calculator and you should get your answer)

OpenStudy (butterflydreamer):